OpenStudy (daedae17):

Find the slope of a line perpendicular to 3x+5y=15 A.-3/5 B.-5/3 C.5/3 D.3/5

OpenStudy (anonymous):

c product of slopes shud be -1

OpenStudy (anonymous):

The new slope is : Calculate the reciprocal of the original slope and then change the sign of the result.\[-\left(\left(\frac{-3}{5}\right)^{-1}\right)=\frac{5}{3} \]